OpenStudy (anonymous):

How do I graph y=x^2 + c when c is -2

OpenStudy (anonymous):

Do you know what the base graph looks like, y=x^2?

OpenStudy (anonymous):

y=ax^2+bx+c

OpenStudy (anonymous):

It wants to know what the affect of c is when c is -2,-1,0,1,2 or as it increases

OpenStudy (anonymous):

Ok, you know y=x^2 is a parabola with vertex at the origin, the c value will shift the graph up or down depending on if its negative or positive

OpenStudy (anonymous):

So the c is what moves it on the y axis?

OpenStudy (anonymous):

Yes it shifts it up or down c units. So if it was -2 it would shift the graph down 2 units

OpenStudy (anonymous):

Oh ok I see.. Thank you!

OpenStudy (anonymous):

So does that mean that the b value determines there placement on the x value?

OpenStudy (anonymous):

What's the b value?

OpenStudy (anonymous):

they give me the same equation but different values for the b (-2 -1 0 1 2 ) to compare with the c value

OpenStudy (anonymous):

Where is the b value in the equation? For the first one you had an example where c was y=x^2+c

OpenStudy (anonymous):

y=ax^2 + bx + c

OpenStudy (anonymous):

Im actually not sure about that one...=p

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

Well thanks anyway lol